Problem 6-20 CVP Applications: Break-Even Analysis; Cost Structure; Target Sales (LO6-1, LO6-3, LOG- 4, LO6-5, LO6-6, LO6-8]Complete this question by entering your answers in the tabs below. Req 1 Reg 2 Req3 Req 4 Req 5 Req 6A Req 6B Compute (a) lasReq 1 Req 2 Req 3 Req 4 Reg 5 Req 6A Req 6B Due to an increase in labor rates, the company estimates that next years variablComplete this question by entering your answers in the tabs below. Req 1 Req 2 Req3 Reg 4 Req 5 Req 6A Req 6B Refer to the daComplete this question by entering your answers in the tabs below. Req 1 Req 2 Req3 Req 4 Req 5 Req 6A Req 6B Refer again toComplete this question by entering your answers in the tabs below. Reg 1 Req 2 Req3 Req 4 Req 5 Req 6A Req 6B Refer to the orComplete this question by entering your answers in the tabs below. Req 1 Req 2 Req3 Req 4 Req 5 Req 6A Req 6B If the new planComplete this question by entering your answers in the tabs below. Req 1 Req 2 Req3 Req 4 Req 5 Req 6A Req 6B Assume the new
Answer #1

As From part 1 to Part 4 are correct, I have only done the remaining parts.

5) New variable cost per unit= ($750000/50000)*60%= $9 per unit

Selling price per unit= $1250000/50000= $25

New fixed costs= $320000*2= $640000

Contribution margin per unit= Selling price per unit-New variable cost per unit

= $25-9= $16 per unit

Contribution margin ratio= Contribution margin*100/Sales

= $16*100/25= 64%

Break even point in balls= Fixed cost/Contribution margin per unit

= $640000/16= 40000 balls

CM ratio 64 %
Unit sales to break even 40000 balls

6-A) Break eve point in balls= (Fixed cost+Target income)/Contribution margin per unit

= ($640000+180000)/16= 51250 balls

Number of balls 51250

B)

Northwood Company
Contribution Income Statement
Sales (50000*$25) $1250000
Variable expenses (50000*$9) -450000
Contribution margin 800000
Fixed expenses -640000
Net operating income $160000
Degree of operating leverage (800000/160000) 5